OpenStudy (ffuentes6596):

Find the product of 2x4(4x2 + 3x + 1). 8x6 + 6x5 + 2x4 8x8 + 3x4 + 2x4 2x4 + 6x5 + 8x6 6x6 + 5x5 + 3x4 (Medals)

OpenStudy (wanell):

2. The correct answer among the choices provided is 8x6 + 6x5 + 2x4.

OpenStudy (justin234):

1. 8x^6 + 6x^5 + 2x^4 2. 8x^3 - 28x^2 + 4x - 6x^2 + 21x - 3 = 8x^3 - 34x^2 + 25x - 3 3. 2x^4 + 4x^3 + 24x + x^3 + 4x^2 + 24x - 6x^2 - 3x - 18 = 2x^4 + 5x^3 - 2x^2 + 45x - 18 4. 2x^2 + 10x - 4x - 20 = 2x^2 - 14x - 20 5. Distributive: x • x - 3 • x - 2 • x + (-2)(-3) = x^2 - 3x - 2x + 6 = x^2 - 5x + 6 so the answer is 8x6 + 6x5 + 2x4.

OpenStudy (justin234):

I explained it and I hope it helps